Describe Gibbs phase rule


1
Expert's answer
2021-06-29T00:31:26-0400

The Gibbs phase rule recognizes the degree of freedom of a multiphase system that is in thermodynamic equilibrium. The law relates the number of intensive independent thermodynamic properties for each phase and the number of phases for a system.

The number of variables which are required to describe a system’s state:

p+f=c+2

Therefore, f=c-p+2

Where,

           p=number of phases

           c=number of components

           f=degrees of freedom

Phase is a region of material that is chemically uniform, physically distinct, and mechanically separated.

Component is minimum number of independent species necessary to define the composition of all phases of the system

Degree of freedom is the number of intensive variable that are independent of each other.